# In-Context Learning Is Provably Bayesian Inference: A Generalization Theory for Meta-Learning

### v1

@@ 1 Introduction
Large language models (LLMs) have moved far beyond classic NLP benchmarks into complex, real-world workflows (Naveed et al., 2024; Zhao et al., 2025) such as code assistants and generators (GitHub, 2025; Team et al., 2025) in software engineering, Med-PaLM 2 (Singhal et al., 2025) in healthcare, text-to-SQL systems (Gao et al., 2024; Shi et al., 2024) in business intelligence, and vision-language-action models (Kim et al., 2024b; Zitkovich et al., 2023) in robotics. In particular, since GPT-3, modern LLMs have demonstrated a striking ability to adapt to new tasks from only a handful of input-output exemplars, without parameter updates (Brown et al., 2020). This phenomenon, known as in-context learning (ICL), appears across diverse datasets and task formats and is at the heart of these workflows (Min et al., 2022; Dong et al., 2024). These deployments share common constraints: inference-time (test-time) prompts are short, and upstream pretraining covers heterogeneous task types. A concrete, finite-sample account of predictive error under these constraints is therefore of key importance to practitioners.
Numerous studies aim to elucidate the behavior of ICL. Wang et al. (2023); Akyürek et al. (2023); von Oswald et al. (2023); Li et al. (2023); Bai et al. (2023); Garg et al. (2022); Mahankali et al. (2024) have empirically or theoretically shown that Transformers can implement canonical estimators and learning procedures in context (e.g., least squares, ridge, and Lasso, gradient-descent steps, model selection), sometimes achieving near-Bayes-optimal performance on linear tasks. Concurrently, Jeon et al. (2024) provide information-theoretic analysis and Kim et al. (2024a) present nonparametric rates for particular architectures and settings, with subsequent progress (Wang et al., 2024; Oko et al., 2024; Nishikawa et al., 2025). A compelling perspective frames ICL as a form of implicit Bayesian inference (Xie et al., 2022; Wang et al., 2023; Panwar et al., 2024; Arora et al., 2025; Reuter et al., 2025; Zhang et al., 2025). Although this viewpoint provides an explanatory framework for ICL’s capabilities, the aforementioned theories have not fully leveraged the theoretical relationship between ICL and Bayes. Hence, they lack a statistical theory that can (i) jointly couple pretraining size $N$ and prompt length $p$ and (ii) accommodate heterogeneous mixtures of task types, the regime in which modern LLMs operate.
We develop a Bayes-centric framework that offers a concrete account of the sources of error and clarifies how they shrink with $p$ and $N$. Specifically, viewing ICL risk as the Bayes risk (e.g., §5.3.1.2 of Murphy, 2022), we treat the Bayes-optimal predictor as the optimal in-context predictor and derive the following orthogonal decomposition under squared loss (Theorem 1):
@@
    Provide non-asymptotic upper bounds that couple the number of pretraining prompts $N$ and their context length $p$ (Theorem 2). For uniform-attention Transformers, we leverage sequential learning theory (Rakhlin et al., 2010), develop optimal transport-based approximation theory, and then obtain
@@
    Here $m$ is the number of learned features in the Transformer, $d_{\mathrm{eff}}$ is the effective dimension, and $\alpha$ is a Hölder exponent. The rate $\propto m/(pN)$ clarifies the dependence on both $p$ and $N$, which earlier theories on ICL (Kim et al., 2024a; Wu et al., 2024; Zhang et al., 2024) have not fully captured. Importantly, the result suggests that Transformers select the optimal meta-algorithm during pretraining.
@@
    Explain in-context error via the test-task difficulty (Theorem 3). In a mixture of task types, the posterior over the task index concentrates exponentially fast with respect to the observed context length, and the irreducible term $R_{\mathrm{PV}}$ is upper bounded by the minimax risk of the test (true) task family. Without assuming specific algorithms (Akyürek et al., 2023; Bai et al., 2023; Zhang et al., 2024), our result implies that even in mixed-task settings the optimal meta-algorithm rapidly converges to the optimal algorithm for the true task at inference time. This finding is consistent with empirical reports (Panwar et al., 2024; Arora et al., 2025), which show that ICL often behaves like Bayesian inference, particularly in task-mixture settings.
@@
    Characterize stability under input-distribution shift (Theorem 4). We demonstrate that under input-distribution shift from pretraining data to inference-time prompt, the Bayes Gap incurs an out-of-distribution (OOD) penalty proportional to the Wasserstein distance between the distributions, while the Posterior Variance is intrinsic to the target domain. Zhang et al. (2024) have noted that ICL is vulnerable to input-distribution shift in some settings, whereas our results specifically show that only the Bayes Gap increases in proportion to the magnitude of the shift.
The paper is organized as follows. Section 2 formalizes the meta-learning prompt model, introduces the Transformer architecture, and states assumptions, followed by a primer on the Bayes-optimal in-context predictor. Section 3 presents the risk decomposition and then analyzes (i) the Bayes Gap (Section 3.1), (ii) the Posterior Variance (Section 3.2), and (iii) OOD stability under input-distribution shift (Section 3.3). Section 4 concludes with limitations and future work. The Appendix contains a list of notation, all technical proofs, auxiliary lemmas, and extended discussions.
@@
(A) ICL as Bayesian inference. ICL has been framed as (implicit) Bayesian inference under structured pretraining. Xie et al. (2022) show that mixtures of hidden Markov model-style documents enable Transformers to perform posterior prediction; Panwar et al. (2024) show that Transformers mimic Bayes across task mixtures. Lin & Lee (2024) reconcile task retrieval versus task learning with a probabilistic pretraining model. Wang et al. (2023) view LLMs as latent-variable predictors enabling principled exemplar selection. Reuter et al. (2025) empirically show full Bayesian posterior inference in-context and Arora et al. (2025) demonstrate Bayesian scaling laws predicting many-shot reemergence of suppressed behaviors. Our results explicitly use Bayesian properties for ICL theory and provide a concrete non-asymptotic validation both in pretraining and at inference time.
(B) ICL as Meta-learning. ICL is widely understood as meta-learning. Transformers implement gradient-descent-style updates within their forward pass, acting as meta-optimizers that perform implicit fine-tuning (von Oswald et al., 2023; Dai et al., 2023). Models can be meta-trained to execute general-purpose in-context algorithms across tasks (Kirsch et al., 2022). From a learning-to-learn perspective, ICL’s expressivity explains few-shot strength while exposing generalization limits (Wu et al., 2025). Beyond single tasks, meta-in-context learning shows recursive adaptation of ICL strategies without parameter updates (Coda-Forno et al., 2023). From this perspective, we theoretically clarify how ICL identifies the task at inference time and solves the true task.
@@ 2.2 Transformer Architecture
We begin by briefly reviewing the standard Transformer architecture. The standard Transformer (Vaswani et al., 2017) processes sequences through self-attention mechanisms: $\text{Attention}(Q,K,V)=\text{softmax}\big(\tfrac{QK^{\top}}{\sqrt{d_{k}}}\big)V$, where queries $Q$, keys $K$, and values $V$ are linear projections of the input embeddings. Each Transformer layer consists of self-attention and a position-wise feed-forward network.
In this work, we adopt a specialized uniform-attention ($Q=K=0$) Transformer architecture. The components of our prompts are generated independently conditional on the task function (Definition 1). Therefore, a permutation-invariant mechanism like uniform attention is sufficient, which motivates our choice of the following architecture. Further justification is provided in Appendix B.
@@ 2.4 Primer on the Bayes-Optimal In-Context Predictor
In this section, we characterize the optimal predictor that minimizes the ICL risk. Since the ICL risk is equivalent to the Bayes risk (e.g., §5.3.1.2 of Murphy, 2022), the theoretically optimal in-context predictor is the Bayes predictor, i.e., the posterior mean of the function value given the context in this setting. We explain this point below.
@@
To minimize the outer expectation, it suffices to minimize the inner conditional expectation for each fixed context $D^{k}$. The minimizer is exactly the definition of the Bayes estimator (Bernardo & Smith, 1994; Robert, 2007) because the inner conditional expectation is the Bayes risk, which is the expected predictive loss $\mathbb{E}_{\bm{x}_{k+1}\sim{\mathcal{P}}_{X}}\left[\ell\left(f(\bm{x}_{k+1}),{M}(P^{k})\right)\right]$ with respect to the Bayes posterior distribution $\mathbb{E}_{I\sim{\mathcal{P}}_{I\mid D^{k}}}[{\mathcal{P}}_{F_{I}\mid D^{k}}]$. Specifically, for the squared error loss, the value $M(P^{k})$ that minimizes the conditional mean squared error, $\mathbb{E}_{I\sim{\mathcal{P}}_{I\mid D^{k}}}\mathbb{E}_{f\sim{\mathcal{P}}_{F_{I}\mid D^{k}}}\mathbb{E}_{\bm{x}_{k+1}\sim{\mathcal{P}}_{X}}\left[\ell\left(f(\bm{x}_{k+1}),{M}(P^{k})\right)\right]$, is the Bayes posterior mean (e.g., Murphy, 2022; Lehmann & Casella, 1998). Thus, the optimal predictor $M_{\mathrm{Bayes}}$ that minimizes the ICL risk is the posterior mean:
@@
This Bayes predictor serves as the theoretical target during pretraining. (See Figure 1.) The Bayes Gap, which we introduce next, measures how well the pretrained model $M_{\hat{\theta}}$ emulates this predictor.
@@
Let $\pi_{i}(D^{k}):=\Pr(I=i\mid D^{k})$ and ${\mathcal{P}}(f\mid D^{k})=\sum_{i=1}^{T}\pi_{i}(D^{k}){\mathcal{P}}_{F_{i}}(f\mid D^{k},I=i)$. We write the Bayes predictor as $M_{\mathrm{Bayes}}(P^{k})=\mathbb{E}_{f\sim{\mathcal{P}}(f\mid D^{k})}[f(\bm{x}_{k+1})]$. Note that, throughout, we work on standard Borel spaces so that regular conditional distributions exist. Accordingly, $\Pr(f\in\cdot\mid D^{k})$ and the quantities $\mathbb{E}[f(\bm{x}_{k+1})\mid D^{k}]$ and $\mathrm{Var}(f(\bm{x}_{k+1})\mid D^{k})$ are well-defined.
@@
For each $k$, we write $\bm{u}_{k}=(\bm{x}_{k},y_{k})\in{\mathcal{U}}$ and $\bm{c}=\bm{x}_{k+1}\in{\mathcal{C}}$ and view the Bayes predictor $M_{\mathrm{Bayes}}(P^{k})$ as $M_{\mathrm{Bayes}}(\bm{u}_{1:k},\bm{c})$ here. Since the posterior ${\mathcal{P}}(f\mid D^{k})$ depends on $D^{k}$ only through the multiset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{k}$, for any permutation $\pi$ of $\{1,\dots,k\}$, $M_{\mathrm{Bayes}}(\bm{u}_{1:k},\bm{c})=M_{\mathrm{Bayes}}\left(\bm{u}_{\pi(1)},\ldots,\bm{u}_{\pi(k)},\bm{c}\right)$. Thus, the Bayes predictor is a symmetric set functional, which justifies using the uniform-attention Transformer to emulate it. See Appendix B for more details.
@@ 3 Risk Analysis of In-Context Learning
This decomposition reveals how each term can be reduced. The Bayes Gap can be controlled through architecture design and pretraining scale ($N,p$). In contrast, the Posterior Variance stems from the inference-time uncertainty of the test task and can be reduced only by increasing the context length $k$ at inference time because $\mathbb{E}[\mathrm{Var}_{f\sim{\mathcal{P}}(f\mid D^{k+1})}(f)]\leq\mathbb{E}[\mathrm{Var}_{f\sim{\mathcal{P}}(f\mid D^{k})}(f)]$ follows from the law of total variance. Therefore, under sufficiently large pretraining, the final error bottleneck is the latter. Also, relative to information-theoretic decompositions under log-loss (Jeon et al., 2024), our identity is exact under squared loss and directly interprets the irreducible term as the Posterior Variance.
@@ 3.1 Bayes Gap: Pretraining Generalization Error and Approximation Error
Proof Idea: Regarding the pretraining generalization error, we handle the $N$ meta-training prompts via conventional learning theory across $j$ (van der Vaart & Wellner, 2023; Shalev-Shwartz & Ben-David, 2014), and the $p$ context examples per prompt via a sequential learning theory across $k$ (Rakhlin et al., 2015; Block et al., 2021). Concerning the approximation error, we build a mollified partition-of-unity (“soft histogram”) over the example domain $\mathcal{U}$} and mean-pool it to encode prompts. Then the Bayes predictor on empirical measures is approximated by a decoder defined via a McShane extension over a discrete 1-Wasserstein metric between histograms (Peyré & Cuturi, 2019), yielding a Lipschitz, piecewise-linear target. Both encoder and decoder are then realized by moderate-size ReLU networks.
The key point of Theorem 2 is that $R_{\mathrm{BG}}$ decomposes into (i) an approximation error $m^{-2\alpha/d_{\mathrm{eff}}}$ stemming from the expressiveness of the Transformer, and (ii) a generalization error $\tilde{O}\!\big(m/(pN)+1/N\big)$ coming from a finite dataset. This decomposition clarifies the respective roles of the two terms. The feature dimension $m$ governs the expressive power of the Transformer, and increasing $m$ allows a smoother approximation of the Bayesian predictor. On the other hand, $p$ represents the amount of information within one task, while $N$ represents the coverage of the meta-distribution. The rate $\propto m/(pN)$ makes explicit the joint effect of $pN$, which earlier non-asymptotic theories on ICL (Kim et al., 2024a; Wu et al., 2024; Zhang et al., 2024) have not fully captured, as they typically considered the effect of $p$ and $N$ separately or focused on only one of them. Many works (e.g., Akyürek et al., 2023; Bai et al., 2023; Zhang et al., 2024) have theoretically and empirically shown that Transformers approximate ridge regression and gradient descent in linear settings. By contrast, we non-asymptotically demonstrate that in more general settings (nonparametric, nonlinear, meta-learning), the optimal meta-algorithm is selected.
@@
The Hölder condition holds, intuitively, if (i) each task function is smooth (e.g., Hölder) with respect to the input, (ii) inputs and responses are effectively bounded (e.g., sub-Gaussian noise), and (iii) Bayesian updates are stable (e.g., distributions of parameters are light-tailed or log-concave), so perturbing any single context point by $O(\delta)$ changes the posterior mean by at most $O(\delta^{\alpha}/k)$ under the prompt metric. These conditions are typically met for mixtures of common task families (e.g., linear regression, basis-function regression, finite convex-dictionary regression). Further discussion is deferred to Appendix D. Moreover, the rate $(pN)^{-\frac{2\alpha}{d_{\mathrm{eff}}+2\alpha}}$ matches the minimax lower bound for estimating, for example, the density of the joint distribution of $(\bm{x}_{i},y_{i})\in{\mathcal{U}}$ under the standard Hölder smoothness assumption (Tsybakov, 2009).

### v2

@@ 1 Introduction
Large language models (LLMs) have moved far beyond classic NLP benchmarks into complex, real-world workflows (naveed2024comprehensiveoverviewlargelanguage; zhao2025surveylargelanguagemodels) such as code assistants and generators (github_copilot; geminiteam2025geminifamilyhighlycapable) in software engineering, Med-PaLM 2 (Singhal2025MedPaLM2) in healthcare, text-to-SQL systems (gao2024text2sql; shi2024survey_tosql) in business intelligence, and vision-language-action models (kim24openvla; pmlr-v229-zitkovich23a) in robotics. In particular, since GPT-3, modern LLMs have demonstrated a striking ability to adapt to new tasks from only a handful of input-output exemplars, without parameter updates (brown2020Language). This phenomenon, known as in-context learning (ICL), appears across diverse datasets and task formats and is at the heart of these workflows (min2022rethinking; dong-etal-2024-survey). These deployments share common constraints: inference-time (test-time) prompts are short, and upstream pretraining covers heterogeneous task types. A concrete, finite-sample account of predictive error under these constraints is therefore of key importance to practitioners.
Numerous studies aim to elucidate the behavior of ICL. wang2023large; akyurek2023what; vonOswald2023transformer; Li2023algorithm; bai2023transformers; garg2022case; mahankali2024one have empirically or theoretically shown that Transformers can implement canonical estimators and learning procedures in context (e.g., least squares, ridge, and Lasso, gradient-descent steps, model selection), sometimes achieving near-Bayes-optimal performance on linear tasks. Concurrently, jeon2024 provide information-theoretic analysis and kim2024transformers present nonparametric rates for particular architectures and settings, with subsequent progress (wang2024incontext; oko2024pretrained; nishikawa2025nonlinear). A compelling perspective frames ICL as a form of implicit Bayesian inference (xie2022an; wang2023large; panwar2024incontext; arora2025bayesian; reuter2025incontext; pmlr-v258-zhang25d). Although this viewpoint provides an explanatory framework for ICL’s capabilities, the aforementioned theories have not fully leveraged the theoretical relationship between ICL and Bayes. Hence, they lack a statistical theory that can (i) jointly couple pretraining size $N$ and prompt length $p$ and (ii) accommodate heterogeneous mixtures of task types, the regime in which modern LLMs operate.
We develop a Bayes-centric framework that offers a concrete account of the sources of error and clarifies how they shrink with $p$ and $N$. Specifically, viewing ICL risk as the Bayes risk (e.g., §5.3.1.2 of Murphy2022intro), we treat the Bayes-optimal predictor as the optimal in-context predictor and derive the following orthogonal decomposition under squared loss (Theorem 1):
@@
    Provide non-asymptotic upper bounds that couple the number of pretraining prompts $N$ and their context length $p$ (Theorem 2). For uniform-attention Transformers, we leverage sequential learning theory (Rakhlin2010online), develop optimal transport-based approximation theory, and then obtain
@@
    Here $m$ is the number of learned features in the Transformer, $d_{\mathrm{eff}}$ is the effective dimension, and $\alpha$ is a Hölder exponent. The rate $\propto m/(pN)$ clarifies the dependence on both $p$ and $N$, which earlier theories on ICL (kim2024transformers; wu2024how; zhang2024trained) have not fully captured. Importantly, the result suggests that Transformers select the optimal meta-algorithm during pretraining.
@@
    Explain in-context error via the test-task difficulty (Theorem 3). In a mixture of task types, the posterior over the task index concentrates exponentially fast with respect to the observed context length, and the irreducible term $R_{\mathrm{PV}}$ is upper bounded by the minimax risk of the test (true) task family. Without assuming specific algorithms (akyurek2023what; bai2023transformers; zhang2024trained), our result implies that even in mixed-task settings the optimal meta-algorithm rapidly converges to the optimal algorithm for the true task at inference time. This finding is consistent with empirical reports (panwar2024incontext; arora2025bayesian), which show that ICL often behaves like Bayesian inference, particularly in task-mixture settings.
@@
    Characterize stability under input-distribution shift (Theorem 4). We demonstrate that under input-distribution shift from pretraining data to inference-time prompt, the Bayes Gap incurs an out-of-distribution (OOD) penalty proportional to the Wasserstein distance between the distributions, while the Posterior Variance is intrinsic to the target domain. zhang2024trained have noted that ICL is vulnerable to input-distribution shift in some settings, whereas our results specifically show that only the Bayes Gap increases in proportion to the magnitude of the shift.
The paper is organized as follows. Section 2 formalizes the meta-learning prompt model, introduces the Transformer architecture, and states assumptions, followed by a primer on the Bayes-optimal in-context predictor. Section 3 presents the risk decomposition and then analyzes (i) the Bayes Gap (Section 3.1), (ii) the Posterior Variance (Section 3.2), and (iii) OOD stability under input-distribution shift (Section 3.3). Section 4 concludes with limitations and future work. The Appendix contains a list of notation, numerical experiments, all technical proofs, auxiliary lemmas, and extended discussions.
@@
(A) ICL as Bayesian inference. ICL has been framed as (implicit) Bayesian inference under structured pretraining. xie2022an show that mixtures of hidden Markov model-style documents enable Transformers to perform posterior prediction; panwar2024incontext show that Transformers mimic Bayes across task mixtures. Lin2024 reconcile task retrieval versus task learning with a probabilistic pretraining model. wang2023large view LLMs as latent-variable predictors enabling principled exemplar selection. reuter2025incontext empirically show full Bayesian posterior inference in-context and arora2025bayesian demonstrate Bayesian scaling laws predicting many-shot reemergence of suppressed behaviors. Our results explicitly use Bayesian properties for ICL theory and provide a concrete non-asymptotic validation both in pretraining and at inference time. Note that, independently and concurrently with our work, ma2025provable have analyzed the adaptivity of ICL from a similar standpoint, but their results are complementary to ours, and neither work encompasses the other.
(B) ICL as Meta-Learning. ICL is widely understood as meta-learning (brown2020Language). Transformers implement gradient-descent-style updates within their forward pass, acting as meta-optimizers that perform implicit fine-tuning (vonOswald2023transformer; dai2022). Models can be meta-trained to execute general-purpose in-context algorithms across tasks (kirsch2022). From a learning-to-learn perspective, ICL’s expressivity explains few-shot strength while exposing generalization limits (wu2025). Beyond single tasks, meta-in-context learning shows recursive adaptation of ICL strategies without parameter updates (codaforno2023). From this perspective, we theoretically clarify how ICL identifies the task at inference time and solves the true task.
@@ 2.2 Transformer Architecture
We begin by briefly reviewing the standard Transformer architecture. The standard Transformer (vaswani2017attention) processes sequences through self-attention mechanisms: $\text{Attention}(Q,K,V)=\text{softmax}\big(\tfrac{QK^{\top}}{\sqrt{d_{k}}}\big)V$, where queries $Q$, keys $K$, and values $V$ are linear projections of the input embeddings. Each Transformer layer consists of self-attention and a position-wise feed-forward network.
In this work, we adopt a specialized uniform-attention ($Q=K=0$) Transformer architecture. The components of our prompts are generated independently conditional on the task function (Definition 1). Therefore, a permutation-invariant mechanism like uniform attention is sufficient, which motivates our choice of the following architecture. Further justification is provided in Appendix C.
@@ 2.4 Primer on the Bayes-Optimal In-Context Predictor
In this section, we characterize the optimal predictor that minimizes the ICL risk. Since the ICL risk is equivalent to the Bayes risk (e.g., §5.3.1.2 of Murphy2022intro), the theoretically optimal in-context predictor is the Bayes predictor, i.e., the posterior mean of the function value given the context in this setting. We explain this point below.
@@
To minimize the outer expectation, it suffices to minimize the inner conditional expectation for each fixed context $D^{k}$. The minimizer is exactly the definition of the Bayes estimator (BernardoSmith1994; robert2007bayesian) because the inner conditional expectation is the Bayes risk, which is the expected predictive loss $\mathbb{E}_{\bm{x}_{k+1}\sim{\mathcal{P}}_{X}}\left[\ell\left(f(\bm{x}_{k+1}),{M}(P^{k})\right)\right]$ with respect to the Bayes posterior distribution $\mathbb{E}_{I\sim{\mathcal{P}}_{I\mid D^{k}}}[{\mathcal{P}}_{F_{I}\mid D^{k}}]$. Specifically, for the squared error loss, the value $M(P^{k})$ that minimizes the conditional mean squared error, $\mathbb{E}_{I\sim{\mathcal{P}}_{I\mid D^{k}}}\mathbb{E}_{f\sim{\mathcal{P}}_{F_{I}\mid D^{k}}}\mathbb{E}_{\bm{x}_{k+1}\sim{\mathcal{P}}_{X}}\left[\ell\left(f(\bm{x}_{k+1}),{M}(P^{k})\right)\right]$, is the Bayes posterior mean (e.g., Murphy2022intro; LehmannCasella1998). Thus, the optimal predictor $M_{\mathrm{Bayes}}$ that minimizes the ICL risk is the posterior mean:
@@
This Bayes predictor serves as the theoretical target during pretraining. (See Figure 1.) Intuitively, the optimal ICL can be viewed as performing implicit prompt learning (li2021prefixtuning; lester-etal-2021-power): given a context $D^{k}$, it infers a task-specific representation ($\mathbb{E}_{I\sim{\mathcal{P}}_{I\mid D^{k}}}[{\mathcal{P}}_{F_{I}\mid D^{k}}]$) which plays a similar role as a learned prompt in prompt learning. The Bayes Gap, which we introduce next, measures how well the pretrained model $M_{\hat{\theta}}$ emulates this predictor.
@@
Let $\pi_{i}(D^{k}):=\Pr(I=i\mid D^{k})$ and ${\mathcal{P}}(f\mid D^{k})=\sum_{i=1}^{T}\pi_{i}(D^{k}){\mathcal{P}}_{F_{i}}(f\mid D^{k},I=i)$. We write the Bayes predictor as $M_{\mathrm{Bayes}}(P^{k})=\mathbb{E}_{f\sim{\mathcal{P}}(f\mid D^{k})}[f(\bm{x}_{k+1})]$ [^1] As the query $\bm{x}_{k+1}$ is drawn independently of $f$ and $D^{k}$, ${\mathcal{P}}(f\mid P^{k})={\mathcal{P}}(f\mid D^{k})$.. Note that, throughout, we work on standard Borel spaces so that regular conditional distributions exist. Accordingly, $\Pr(f\in\cdot\mid D^{k})$ and the quantities $\mathbb{E}[f(\bm{x}_{k+1})\mid D^{k}]$ and $\mathrm{Var}(f(\bm{x}_{k+1})\mid D^{k})$ are well-defined.
@@
For each $k$, we write $\bm{u}_{k}=(\bm{x}_{k},y_{k})\in{\mathcal{U}}$ and $\bm{c}=\bm{x}_{k+1}\in{\mathcal{C}}$ and view the Bayes predictor $M_{\mathrm{Bayes}}(P^{k})$ as $M_{\mathrm{Bayes}}(\bm{u}_{1:k},\bm{c})$ here. Since the posterior ${\mathcal{P}}(f\mid D^{k})$ depends on $D^{k}$ only through the multiset $\{(\bm{x}_{i},y_{i})\}_{i=1}^{k}$, for any permutation $\pi$ of $\{1,\dots,k\}$, $M_{\mathrm{Bayes}}(\bm{u}_{1:k},\bm{c})=M_{\mathrm{Bayes}}\left(\bm{u}_{\pi(1)},\ldots,\bm{u}_{\pi(k)},\bm{c}\right)$. Thus, the Bayes predictor is a symmetric set functional, which justifies using the uniform-attention Transformer to emulate it. See Appendix C for more details.
@@ 3 Risk Analysis of In-Context Learning
This decomposition reveals how each term can be reduced. The Bayes Gap is the aggregate imperfections of pretraining, controlled through architecture design and pretraining scale ($N,p$). In contrast, the Posterior Variance stems from the inference-time uncertainty of the test task and can be reduced only by increasing the context length $k$ at inference time because $\mathbb{E}[\mathrm{Var}_{f\sim{\mathcal{P}}(f\mid D^{k+1})}(f)]\leq\mathbb{E}[\mathrm{Var}_{f\sim{\mathcal{P}}(f\mid D^{k})}(f)]$ follows from the law of total variance. Therefore, under sufficiently large pretraining, the final error bottleneck is the latter. Also, relative to information-theoretic decompositions under log-loss (jeon2024), our identity is exact under squared loss and directly interprets the irreducible term as the Posterior Variance. Note that this decomposition holds for any measurable and bounded predictor, independent of the specific form of the model. Also, the analysis can be extended to a broader class of losses that admit an analogous decomposition with the Bayes estimator, such as Bregman-type losses (adlam2022biasvariancebregman; pfau2025generalizedbvd).
@@ 3.1 Bayes Gap: Pretraining Generalization Error and Approximation Error
Proof Idea: Regarding the pretraining generalization error, we handle the $N$ meta-training prompts via conventional learning theory across $j$ (vaart_wellner_weak_2023; Shalev-Shwartz_Ben-David_2014), and the $p$ context examples per prompt via a sequential learning theory across $k$ (Rakhlin2015; BlockDaganRakhlin2021). Concerning the approximation error, we build a mollified partition-of-unity (“soft histogram”) over the example domain $\mathcal{U}$} and mean-pool it to encode prompts. Then the Bayes predictor on empirical measures is approximated by a decoder defined via a McShane extension over a discrete 1-Wasserstein metric between histograms (peyré2020computational), yielding a Lipschitz, piecewise-linear target. Both encoder and decoder are then realized by moderate-size ReLU networks. As these proof ideas do not depend on a specific Bayesian formulation, the result holds under milder data assumptions compared to prior Bayesian analyses (xie2022an; pmlr-v258-zhang25d).
The key point of Theorem 2 is that $R_{\mathrm{BG}}$ decomposes into (i) an approximation error $m^{-2\alpha/d_{\mathrm{eff}}}$ stemming from the expressiveness of the Transformer, and (ii) a generalization error $\tilde{O}\!\big(m/(pN)+1/N\big)$ coming from a finite dataset. This decomposition clarifies the respective roles of the two terms. The feature dimension $m$ governs the expressive power of the Transformer, and increasing $m$ allows a smoother approximation of the Bayesian predictor. The depth and number of parameters of the Transformer affect Theorem 2 only implicitly, through the feature dimension $m$ and the Lipschitz constants of the encoder and decoder (Lemmas 4-5). On the other hand, $p$ represents the amount of information within one task, while $N$ represents the coverage of the meta-distribution. The rate $\propto m/(pN)$ makes explicit the joint effect of $pN$, which earlier non-asymptotic theories on ICL (kim2024transformers; wu2024how; zhang2024trained) have not fully captured, as they typically considered the effect of $p$ and $N$ separately or focused on only one of them. Many works (e.g., akyurek2023what; bai2023transformers; zhang2024trained) have theoretically and empirically shown that Transformers approximate ridge regression and gradient descent in linear settings. By contrast, we non-asymptotically demonstrate that in more general settings (nonparametric, nonlinear, meta-learning), the optimal meta-algorithm is selected.
@@
The Hölder condition holds, intuitively, if (i) each task function is smooth (e.g., Hölder) with respect to the input, (ii) inputs and responses are effectively bounded (e.g., sub-Gaussian noise), and (iii) Bayesian updates are stable (e.g., distributions of parameters are light-tailed or log-concave), so perturbing any single context point by $O(\delta)$ changes the posterior mean by at most $O(\delta^{\alpha}/k)$ under the prompt metric. These conditions are typically met for mixtures of common task families (e.g., linear regression, basis-function regression, finite convex-dictionary regression). Further discussion is deferred to Appendix E. Moreover, the rate $(pN)^{-\frac{2\alpha}{d_{\mathrm{eff}}+2\alpha}}$ matches the minimax lower bound for estimating, for example, the density of the joint distribution of $(\bm{x}_{i},y_{i})\in{\mathcal{U}}$ under the standard Hölder smoothness assumption (Tsybakov2009).